BraydenZinsli:

Which of the purchases would be the least expensive? Purchase Amount of sale Sales tax rate A $150 5% B $145 7% C $140 8% D $135 9% purchase A purchase B purchase C purchase D

Hoodmemes:

Ok first u need to find price of each purchase. (5%) (Divided) by (100) \[5\div100=0.05\] Now (Multiply) (0.05) by (150) \[0.05\times150=7.5\] Last (Add) (7.5) to (150) \[150+7.5=157.5\] (157.5) is also know as (157.50) Option A: (157.50) Lets (Divide) (7%) by (100) \[7\div100=0.07\] Now (Multiply) (0.07) by (145) \[0.07\times145=10.15\] Lastly (Add) (10.15) and (145) \[145 + 10.15 = 155.15\] Option B: (155.15) Now (Divide) (8) by (100) \[8\div100=0.08\] Lets (Multiply) (0.08) by (140) \[0.08\times140=11.2\] Finally lets (Add) (11.2) and (140) \[140+11.2=151.2\] (151.2) is also known as (151.20) Option C: (151.20) First u wanna (Divide) (9%) by (100) \[9\div100=0.09\] Lets (Multiply) (0.09) by (135) \[0.09\times135=12.15\] Lastly (Add) (12.15) to (135) \[135+12.15=147.15\] Option D: (147.15) _______________________________________________________________________ Option A: (157.50) Option B: (155.15) Option C: (151.20) Option D: (147.15) As we can see Option D is the (Lowest) price out of all of the Options. ________________________________________________________________________ Option D is the correct,least expensive purchase.